# A hidden Markov model for statistical arbitrage in international crude oil futures markets

### v2

@@ 2.2. The filter-based Expectation Maximization algorithm
    As far as the initialization of the algorithm is concerned, the initial state of the Markov chain $\mathbf{X}_{0}$ also determines $O^{i}_{0}$. In our analysis we set $\hat{\mathbf{X}}_{0}=\mathbf{e}_{1}$ by convention. This implies $\hat{O}^{1}_{0}=1$ and $\hat{O}^{i}_{0}=0$ for $i=2,\dots,N$. With respect to the initial guess for $\boldsymbol{\Pi}$, if $N=2$, we set $\hat{\pi}_{11}^{(0)}=0.6$ and $\hat{\pi}_{22}^{(0)}=0.5$ since the perfectly symmetric initial guess $\hat{\pi}_{11}^{(0)}=\hat{\pi}_{22}^{(0)}=0.5$ tends to create numerical instabilities as there is no difference across states in the transition matrix. If $N=3$, we set $\hat{\pi}_{11}^{(0)}=0.5,\hat{\pi}_{12}^{(0)}=0.25$, $\hat{\pi}_{21}^{(0)}=0.3,\hat{\pi}_{22}^{(0)}=0.4$ and $\hat{\pi}_{31}^{(0)}=0.2,\hat{\pi}_{33}^{(0)}=0.6$. The initial guesses for $\boldsymbol{\gamma},\boldsymbol{\alpha}$ and $\boldsymbol{\eta}$ are derived from the data. More specifically, if the Markov chain $\mathbf{X}$ was constant, then the process $y$ would follow a simple AR(1) process with constant parameters satisfying
    (2.9) $y_{t+1}=\gamma+\alpha y_{t}+\varepsilon_{t+1},$
@@
    where $\varepsilon=(\varepsilon_{t})_{t\in\mathbb{N}}$ is a Gaussian white noise process with variance $\eta^{2}$. In this case, the parameters $\gamma$, $\alpha$, $\eta$ could be estimated by ordinary least squares. In our empirical analysis we do so using the first $20$ datapoints of $y$, corresponding to one month of daily observations (this is twice the batch parameter $m$ introduced below and that we set equal to $10$). For $N=1$ we therefore set $\hat{\gamma}^{(0)}=\hat{\gamma}_{\rm OLS}$, $\hat{\alpha}^{(0)}=\hat{\alpha}_{\rm OLS}$ and $\hat{\eta}^{(0)}=\hat{\eta}_{\rm OLS}$, where $\hat{\gamma}_{\rm OLS}$, $\hat{\alpha}_{\rm OLS}$ and $\hat{\eta}_{\rm OLS}$ denote the ordinary least square estimates of the parameters of (2.9). For $N=2$ we let the initial guess for the two states for all the parameters be equally spaced with respect to the OLS estimates determined as above. Hence, we set $\hat{\gamma}_{1}^{(0)}=1.3\hat{\gamma}_{\rm OLS}$ and $\hat{\gamma}_{2}^{(0)}=0.7\hat{\gamma}_{\rm OLS}$, and analogously for the other two parameters. For $N=3$ we follow the same reasoning and set $\hat{\gamma}_{1}^{(0)}=1.3\hat{\gamma}_{\rm OLS}$, $\hat{\gamma}_{2}^{(0)}=\hat{\gamma}_{\rm OLS}$ and $\hat{\gamma}_{3}^{(0)}=0.7\hat{\gamma}_{\rm OLS}$, and analogously for the remaining parameters. We point out that the EM algorithm is robust with respect to the specification of the initial guesses.
@@
    Quantities like $\hat{T}_{t}^{i}(y_{t-1}^{2})$, appearing at the denominator in $\hat{\alpha}_{i}^{(t+1)}$ and $\hat{\eta}_{i}^{(t+1)}$ in (2.8), might take very small values (especially when the $y$’s are close to zero) and this might induce numerical instabilities. The same issue can arise when computing the diagonal elements of $\mathbf{D}$ if the $\hat{\eta}_{i}^{(t+1)}$’s are too close to zero. If not controlled for, these instabilities propagate, preventing the algorithm to converge. Therefore, for all denominators in (2.8) we check that the new estimate of the quantity of interest is not smaller/larger than ten times the previous estimate and, in case, we truncate the new estimate to that level.
@@
    The two previous issues can be mitigated by adopting a batchwise approach, as pointed out in [Grimm et al., 2020], updating the parameters in (2.8) not at every time point $t$ as the filters in (2.7) but rather every $m$ steps. This technique stabilizes the estimates of the filtered quantities in (2.7) and, as a consequence, the parameter estimates in (2.8). In our analysis we set $m=10$, which in practice amounts to update the model parameters every two weeks.
@@ 3. Statistical arbitrage strategies
In this section we introduce several statistical arbitrage strategies that shall be tested on data (Section 3.1) and describe the performance measures used to assess their profitability (Section 3.2).
3.1. Statistical arbitrage strategies
@@
As explained in [Vidyamurthy, 2004, Chapter 8], the basic intuition behind pairs trading is that whenever the spread deviates “sufficiently” from its equilibrium value, the investor should open a trading position, appropriately investing in the underlying assets. The position is then closed when this deviation corrects itself and the spread reverts to its equilibrium value. The precise specification of what we mean above by “sufficiently” determines a specific trading strategy, which can be summarized by a simple rule based on opening and closing signals.
@@
Regardless of the specification of opening/closing signals, we assume that the trading position is determined by the cointegration vector $\lambda=(\lambda^{0},\lambda^{B},\lambda^{S},\lambda^{W})$ in (2.1), where $\lambda^{0}$ determines the borrowing/lending of money from a riskless money market account (earning zero interest rate, for simplicity of presentation). Abstracting from transaction costs, if an opening signal is received at time $t$ when $S_{t}>0$ (resp. $S_{t}<0$), the investor has to go short (resp. long) on portfolio $\lambda$. This generates an inflow of money at time $t$ equal to $|S_{t}|$. Closing this position when the spread reverts back to zero will deliver no cashflow. Therefore, these strategies are designed to deliver positive payoffs at their opening, with zero cashflows when they are closed, similarly to arbitrage opportunities of the second kind (see, e.g., [Ingersoll, 1987]). However, since it is not known ex ante if and when $S_{t}$ will revert back to zero, these strategies are not pure arbitrage opportunities but rather statistical arbitrage opportunities (see [Bondarenko, 2003] and [Rein et al., 2021] for a mathematical formalization of the concept of statistical arbitrage).
In our analysis we consider five different statistical arbitrage strategies. The first one is a simple benchmark and requires no explicit modelling of the spread process. The second and the fourth strategies rely only on sample estimates of the moments of the spread. On the contrary, the third and the fifth strategies exploit our stochastic model for the spread, as they are based on the one-step ahead forecast of the spread and on its forecast interval. While the timing of opening signals differs across the five different rules, the closing one is assumed to be the same for simplicity of comparison. In particular, if there was an open position at time $t-1$, the position is closed at time $t$ if the spread $S_{t}$ changes sign (in discrete time, this corresponds to the first passage of the spread at zero, coherently with the above description).
@@
Formally, assuming there is no open position at $t-1$, a position is opened at $t$ if $S_{t}\neq 0$.
@@
Formally, assuming there is no open position at $t-1$, a position is opened at time $t$ if $S_{t}\notin\left(\hat{\mu}_{t-1-n:t-1}\pm q_{\alpha}\hat{\sigma}_{t-1-n:t-1}\right)$, where $\hat{\mu}_{n_{1}:n_{2}}$ (resp. $\hat{\sigma}_{n_{1}:n_{2}}$) is the sample mean (resp. standard deviation) estimated using the datapoints from $n_{1}$ to $n_{2}$ and $q_{\alpha}$ is the $\alpha$-quantile of the standard normal distribution. In our empirical analysis we use the $0.975^{\text{th}}$ quantile. This strategy is equivalent to the one based on the $z$-score (considered for instance in [Avellaneda and Lee, 2010]), defined as $z_{t}:=\frac{S_{t}-\hat{\mu}_{t-1-n:t-1}}{\hat{\sigma}_{t-1-n:t-1}}$, that prescribes to open a position at $t$ if $z_{t}\notin(-q_{\alpha},+q_{\alpha})$.
Strategy 3: prediction interval (PredI). By exploiting our stochastic model of Section 2, we can improve on the previous strategy by adopting a forward-looking approach, inspired by [Elliott et al., 2005]. According to the PredI rule, the interval the investor should consider is not the backward-looking confidence interval introduced in strategy ProbI, but rather the prediction interval where the moments are computed according to the stochastic model of Section 2.1.
Formally, assuming there is no open position at $t-1$, a position is opened at time $t$ if $S_{t}\notin(\mathbb{E}[S_{t}|\mathcal{F}^{y}_{t-1}]\pm q_{\alpha}\sqrt{\mathbb{V}ar[S_{t}|\mathcal{F}^{y}_{t-1}]})$, where $\mathbb{E}[S_{t}|\mathcal{F}^{y}_{t-1}]$ is the one step-ahead forecast of $S_{t}$ and $\mathbb{V}ar[S_{t}|\mathcal{F}^{y}_{t-1}]$ its variance, both computed by relying on the filtering algorithm described in Section 2.2 on the basis of the information generated by spread process $y$ itself. More specifically,
$\mathbb{E}[S_{t}|\mathcal{F}^{y}_{t-1}]=\hat{\gamma}^{(t-1)}(\hat{\mathbf{X}}_{t-1})+\hat{\alpha}^{(t-1)}(\hat{\mathbf{X}}_{t-1})S_{t-1}\qquad\text{and}\qquad\mathbb{V}ar[S_{t}|\mathcal{F}^{y}_{t-1}]=(\hat{\eta}^{(t-1)}(\hat{\mathbf{X}}_{t-1}))^{2}.$
Observe that these quantities depend on the filtered estimate of the latent Markov chain X. Similarly to the ProbI strategy described above, in this strategy we use the $0.975^{\text{th}}$ quantile.
@@
Formally, assuming there is no open position at $t-1$, a position is opened at $t$ if $x_{t}\notin\left(q_{\underline{\alpha},x},q_{\overline{\alpha},x}\right)$ with $\underline{\alpha}<\overline{\alpha}$ and where $q_{\alpha,x}$ is the empirical quantile of order $\alpha$ of the spread increment $x$. In our analysis we consider the $\underline{\alpha}=0.025^{\text{th}}$ and the $\overline{\alpha}=0.975^{\text{th}}$ quantiles.
Strategy 5: predicted increment (PI). By exploiting the stochastic model of Section 2, we can improve on strategy RI described above. More specifically, the investor can compute the series $\hat{x}=(\hat{x}_{t})_{t\in\mathbb{N}}$ of predicted spread increments, with $\hat{x}_{t}:=\mathbb{E}[S_{t+1}|\mathcal{F}^{y}_{t}]/S_{t}-1$, and trade whenever the predicted increment lies outside a given interval.
Formally, assuming there is no open position at $t-1$, a position is opened at time $t$ if $\hat{x}_{t}\notin\left(q_{\underline{\alpha},\hat{x}},q_{\overline{\alpha},\hat{x}}\right)$. In our analysis we consider the $\underline{\alpha}=0.025^{\text{th}}$ and the $\overline{\alpha}=0.975^{\text{th}}$ quantiles.
Finally, we shall also consider two passive strategies that do not incur into transaction costs: a buy-and-hold position on the S&P500 index (S&P strategy), which will be used as a generic benchmark, and a buy-and-hold position on the Invesco DB Oil Fund [^4] See https://www.invesco.com/us/financial-products/etfs/product-detail?audienceType=Investor&ticker=DBO. Numerical experiments show that different alternatives, like the United States Oil Fund ETF or the WisomTree Crude Oil ETF, are characterized by almost identical performances. (ETF strategy), which is one of the most liquid ETFs for crude oil futures markets and represents a simple alternative for an investor willing to invest in the crude oil futures market.

### v3

@@ 2.2. The filter-based Expectation Maximization algorithm
    Concerning the initialization of the algorithm, the initial state of the Markov chain $\mathbf{X}_{0}$ also determines $O^{i}_{0}$. In our analysis we set $\hat{\mathbf{X}}_{0}=\mathbf{e}_{1}$ by convention. This implies $\hat{O}^{1}_{0}=1$ and $\hat{O}^{i}_{0}=0$, for $i=2,\dots,N$. With respect to the initial guess for $\boldsymbol{\Pi}$, if $N=2$, we set $\hat{\pi}_{11}^{(0)}=0.6$ and $\hat{\pi}_{22}^{(0)}=0.5$ since the perfectly symmetric initial guess $\hat{\pi}_{11}^{(0)}=\hat{\pi}_{22}^{(0)}=0.5$ tends to create numerical instabilities as there is no difference across states in the transition matrix. If $N=3$, we set $\hat{\pi}_{11}^{(0)}=0.5,\hat{\pi}_{12}^{(0)}=0.25$, $\hat{\pi}_{21}^{(0)}=0.3,\hat{\pi}_{22}^{(0)}=0.4$ and $\hat{\pi}_{31}^{(0)}=0.2,\hat{\pi}_{33}^{(0)}=0.6$.
  2. (2)
    The initial guesses for the parameters $\boldsymbol{\gamma},\boldsymbol{\alpha}$, $\boldsymbol{\eta}$ are derived from the data. More specifically, in the absence of regime switching, the spread process would follow a simple AR(1) process with constant parameters satisfying
    (2.8) $S_{t+1}=\gamma+\alpha S_{t}+\varepsilon_{t+1},\qquad\text{ for all }t\in\mathbb{N},$
@@
    where $\varepsilon=(\varepsilon_{t})_{t\in\mathbb{N}}$ is a Gaussian white noise process with variance $\eta^{2}$. In this case, the parameters $\gamma$, $\alpha$, $\eta$ can be estimated by ordinary least squares (OLS). In our empirical analysis we do so using the first $20$ datapoints of $S$, corresponding to one month of daily observations (twice the batch parameter $m$ introduced below). For $N=1$, we set $\hat{\gamma}^{(0)}=\hat{\gamma}_{\rm OLS}$, $\hat{\alpha}^{(0)}=\hat{\alpha}_{\rm OLS}$ and $\hat{\eta}^{(0)}=\hat{\eta}_{\rm OLS}$, where $\hat{\gamma}_{\rm OLS}$, $\hat{\alpha}_{\rm OLS}$ and $\hat{\eta}_{\rm OLS}$ denote the OLS estimates of the parameters of (2.8). For $N=2$, we let the initial guesses of all parameters for the two states be equally spaced with respect to the OLS estimates determined as above. Hence, we set $\hat{\gamma}_{1}^{(0)}=1.3\hat{\gamma}_{\rm OLS}$ and $\hat{\gamma}_{2}^{(0)}=0.7\hat{\gamma}_{\rm OLS}$, and analogously for the other parameters. For $N=3$, we follow the same reasoning and set $\hat{\gamma}_{1}^{(0)}=1.3\hat{\gamma}_{\rm OLS}$, $\hat{\gamma}_{2}^{(0)}=\hat{\gamma}_{\rm OLS}$ and $\hat{\gamma}_{3}^{(0)}=0.7\hat{\gamma}_{\rm OLS}$, and similarly for the remaining parameters.
  3. (3)
    Due to the recursive and online nature of the filter-based EM algorithm, the specification of initial parameter values constitutes the only non-data-driven input of the algorithm. This aspect is particularly relevant in regime-switching models, whose likelihood function is generally non-concave and may admit multiple local maxima (see, e.g., [McLachlan and Krishnan, 2008, Chapter 3]). To assess the stability of our results with respect to the initialization and address potential concerns related to convergence to local optima, we perform a systematic set of robustness checks by varying the initial values of all regime-dependent parameters and transition probabilities. As documented in Remarks 4.1 and 4.3, while the resulting point estimates exhibit mild sensitivity to different initializations, the performance of the strategies derived from the model remains stable. This indicates that the results reported in Section 4 are robust with respect to the initialization and are not affected by convergence to a particular local maximum of the likelihood.
@@
    Quantities like $\hat{T}_{t}^{i}(S_{t-1}^{2})$, appearing at the denominator in $\hat{\alpha}_{i}^{(t+1)}$ and $\hat{\eta}_{i}^{(t+1)}$ in (2.7), might take very small values (especially when the $S$’s are close to zero) and this might induce numerical instabilities. The same issue can arise when computing the diagonal elements of $\mathbf{D}$ if the $\hat{\eta}_{i}^{(t+1)}$’s are too close to zero. If not controlled for, these instabilities propagate, preventing the algorithm to converge. Therefore, for all denominators in (2.7) we check that the new estimate of the quantity of interest is not smaller/larger than ten times the previous estimate and, in case, we truncate the new estimate to that level.
@@
    The two previous issues can be mitigated by adopting a batchwise approach, as pointed out in [Grimm et al., 2020], updating the parameters in (2.7) not at every time point $t$ as the filters in (2.6) but rather every $m$ steps. This technique stabilizes the estimates of the filtered quantities in (2.6) and, as a consequence, the parameter estimates in (2.7). In our analysis we set $m=10$, which amounts to updating the model parameters every two weeks.
@@ 3. Statistical arbitrage strategies
In this section we introduce several statistical arbitrage strategies that shall be tested (Section 3.1) and describe the performance measures used to assess their profitability (Section 3.2).
3.1. Construction of statistical arbitrage strategies
@@
As explained in [Vidyamurthy, 2004, Chapter 8], the basic intuition behind pairs trading is that whenever the spread deviates “sufficiently” from its equilibrium value, the investor should open a trading position, appropriately investing in the underlying assets. The position is then closed when this deviation corrects itself and the spread reverts to its equilibrium value. The precise specification of what we mean above by “sufficiently” determines a specific trading strategy, which can be summarized by a simple rule based on opening and closing signals, as explained in the following.
We consider self-financing trading strategies investing in futures contracts. Any cash not engaged in futures positions is implicitly invested in a riskless money market account, earning zero interest rate for simplicity of presentation. When no trading signal is active, the strategy holds no futures positions and therefore generates zero returns. Regardless of the specification of opening/closing signals, we assume that the trading position is determined by the cointegration vector $\lambda=(\lambda^{0},\lambda^{B},\lambda^{S},\lambda^{W})$ in (2.1), where $\lambda^{0}$ represents the cash component. Abstracting from transaction costs, if an opening signal is received at time $t$ when $S_{t}>0$ (resp. $S_{t}<0$), the investor has to go short (resp. long) on portfolio $\lambda$. This generates an inflow of money at time $t$ equal to $|S_{t}|$. A position opened at time t1t-1 is closed at time $t$ if the spread $S_{t}$ changes sign. The cointegration vector $\lambda$ determines the relative composition of the portfolio. Since cointegration vectors are defined only up to scale, a normalization is required to define the trading strategies in a a meaningful and economically interpretable way. In our framework, the scale of the investment at time $t$ is set by normalizing the portfolio $\lambda$ by the gross notional exposure, defined as $G_{t}:=|\lambda_{0}|+\sum_{k=B,S,W}|\lambda_{k}|F^{k}_{t}$, for $t\in\mathbb{N}$. This quantity represents the total market value of the long and short futures positions implied by the cointegration vector at prevailing prices. This normalization will yield returns per unit of gross exposure and ensure that performance measures are invariant to rescaling the cointegration vector (see Section 3.2).
@@
The above trading rule corresponds to the first crossing of the spread through zero in discrete-time, consistent with standard practice in pairs trading. Closing the position as soon as the spread crosses zero will deliver either no cashflow or a small positive one. Therefore, these strategies are designed to deliver positive payoffs at their opening, with essentially zero cashflows when they are closed, similarly to arbitrage opportunities of the second kind (see, e.g., [Ingersoll, 1987]). However, since it is not known ex ante if and when $S_{t}$ will revert back to zero, these strategies are not pure arbitrage opportunities but only statistical arbitrage opportunities (see [Bondarenko, 2003] and [Rein et al., 2021] for a formalization of statistical arbitrage).
Remark 3.1.
We emphasize that, by design, our strategies are not zero-net investment strategies, which are a well-established approach in pairs trading. Instead, they can be described as “negative-cost” investment strategies: they generate a positive cash inflow at initiation and require no cash outflow at closure. Nevertheless, like traditional zero-net investment strategies, they do not require initial capital and are, in principle, scalable. Importantly, consistent with this feature, the returns reported in Subsection 3.2 should be interpreted as excess returns.
In our analysis we consider five different statistical arbitrage strategies. Strategy 1 serves as a simple benchmark and requires no explicit modelling of the spread process. Strategies 2 and 4 are backward-looking in nature and are based on sample estimates of the moments of the spread. By contrast, Strategies 3 and 5 exploit the stochastic model introduced in Section 2, as they are based on one-step-ahead forecast of the spread and on the corresponding forecast interval.
With the exception of the benchmark strategy, all strategies are designed to time both the opening and the closing of trades. These timing decisions are driven by crossings of the observed spread process with a set of bands, defined either as probability intervals or as forecast intervals, depending on the specific strategy. The width of these bands is determined by the quantiles of the relevant distributions at a common confidence level $\alpha$, which we refer to as the “bandwidth level”. The bandwidth level $\alpha$ controls the aggressiveness of the strategies by regulating the frequency of trading signals. Larger values of $\alpha$ yield narrower bands and, therefore, more frequent trading. This raises the gross profit potential by exploiting a larger number of spread deviations, but also amplifies cumulative transaction costs and the exposure to short-lived noisy signals. On the contrary, smaller values of $\alpha$ yield wider bands and more conservative trading strategies, reducing transaction costs at the expense of potentially foregone trading opportunities.
In Subsection 4.2, we examine the sensitivity of strategy performance to different choices of $\alpha$, showing that some strategies exhibit robust performance across a wide range of values. For all further experiments, we consider $\alpha=0.20$, which represents an intermediate choice within the range commonly adopted in the literature, which spans from $\alpha=0.30$ (bands of approximately one standard deviation) to $\alpha=0.05$ (bands of approximately two standard deviations).
@@
Formally, assuming there is no open position at t1t-1, a position is opened at $t$ if $S_{t}\neq 0$.
@@
Formally, assuming there is no open position at t1t-1, a position is opened at time $t$ if $S_{t}\notin(\hat{\mu}_{t-1-n:t-1}\pm|q_{\frac{\alpha}{2}}|\hat{\sigma}_{t-1-n:t-1})$, where $\hat{\mu}_{n_{1}:n_{2}}$ (resp. $\hat{\sigma}_{n_{1}:n_{2}}$) is the sample mean (resp. standard deviation) estimated using the datapoints from $n_{1}$ to $n_{2}$, $q_{\frac{\alpha}{2}}$ is the $\frac{\alpha}{2}$-quantile of the standard normal distribution and $\alpha$ is the previously described bandwidth level. This strategy is equivalent to the one based on the $z$-score (considered for instance in [Avellaneda and Lee, 2010]), defined as $z_{t}:=\frac{S_{t}-\hat{\mu}_{t-1-n:t-1}}{\hat{\sigma}_{t-1-n:t-1}}$, that prescribes to open a position at $t$ if $z_{t}\notin(-q_{\frac{\alpha}{2}},+q_{\frac{\alpha}{2}})$.
Strategy 3: prediction interval (PredI). By exploiting our stochastic model of Section 2, we can improve on the previous strategy by adopting a forward-looking approach, inspired by [Elliott et al., 2005]. According to the PredI rule, the interval the investor should consider is not the backward-looking probability interval introduced in strategy ProbI, but rather the prediction interval where the moments are computed according to the stochastic model of Section 2.1.
Formally, assuming there is no open position at t1t-1, a position is opened at time $t$ if $S_{t}\notin(\mathbb{E}[S_{t}|\mathcal{F}^{S}_{t-1}]\pm|q_{\frac{\alpha}{2}}|\sqrt{\mathbb{V}ar[S_{t}|\mathcal{F}^{S}_{t-1}]})$, where $\mathbb{E}[S_{t}|\mathcal{F}^{S}_{t-1}]$ is the one step-ahead forecast of $S_{t}$ and $\mathbb{V}ar[S_{t}|\mathcal{F}^{S}_{t-1}]$ its variance, both computed by relying on the filtering algorithm described in Section 2.2 on the basis of the information generated by spread process $S$ itself. More specifically,
$\mathbb{E}[S_{t}|\mathcal{F}^{S}_{t-1}]=\hat{\gamma}^{(t-1)}(\hat{\mathbf{X}}_{t-1})+\hat{\alpha}^{(t-1)}(\hat{\mathbf{X}}_{t-1})S_{t-1}\qquad\text{and}\qquad\mathbb{V}ar[S_{t}|\mathcal{F}^{S}_{t-1}]=(\hat{\eta}^{(t-1)}(\hat{\mathbf{X}}_{t-1}))^{2}.$
Observe that these quantities depend on the filtered estimate of the latent Markov chain X.
@@
Formally, assuming there is no open position at t1t-1, a position is opened at time $t$ if $x_{t}\notin(-|q_{\frac{\alpha}{2},x}|,|q_{\frac{\alpha}{2},x}|)$ where $q_{\frac{\alpha}{2},x}$ is the empirical quantile of order $\frac{\alpha}{2}$ of the spread increment $x$.
Strategy 5: predicted increment (PI). By exploiting the stochastic model of Section 2, we can improve on strategy RI described above. More specifically, the investor can compute the series $\hat{x}=(\hat{x}_{t})_{t\in\mathbb{N}}$ of predicted spread increments, with $\hat{x}_{t}:=\mathbb{E}[S_{t+1}|\mathcal{F}^{S}_{t}]/S_{t}-1$, and trade whenever the predicted increment lies outside a given interval.
Formally, assuming there is no open position at t1t-1, a position is opened at time $t$ if $\hat{x}_{t}\notin(-|q_{\frac{\alpha}{2},x}|,|q_{\frac{\alpha}{2},x}|)$.
Finally, we shall also consider two passive strategies that do not incur into transaction costs: a buy-and-hold position on the S&P500 index (Ex. S&P), which will be used as a generic benchmark, and a buy-and-hold position on the Invesco DB Oil Fund [^4] See https://www.invesco.com/us/financial-products/etfs/product-detail?audienceType=Investor&ticker=DBO. Numerical experiments show that different alternatives, like the United States Oil Fund ETF or the WisomTree Crude Oil ETF, are characterized by almost identical performances. (Ex. ETF), which is one of the most liquid ETFs for crude oil futures markets and represents a simple alternative for an investor willing to invest in the crude oil futures market. For sake of comparison, the returns of both these strategies are computed in excess of the prevailing risk-free rate.